OpenStudy (anonymous):

I have the answer can someone please check me ... Find the derivative m(t)=(5)cos^2t

OpenStudy (anonymous):

\[5\cos^2(t)\]?

OpenStudy (anonymous):

The equation for the question is \[m(t)=\frac{ 5 }{ \cos^2t }\]

OpenStudy (anonymous):

\[5\sec^2(t)\]?

OpenStudy (anonymous):

it will be \(10\sec(t)\times \frac{d}{dx}\sec(t)\) or \[10\sec(t)\sec(t)\tan(t)\] or even \[10\sec^2(t)\tan(t)\]

OpenStudy (anonymous):

could the answer be\[m'(t)=\frac{ 5 }{ 2sint cost }\]

OpenStudy (anonymous):

no

OpenStudy (anonymous):

it is easiest to turn in to \(5\sec^2(t)\) but if for some reason you want to work with \(\frac{5}{\cos^2(t)}\) you have to either use the quotient rule or the rule for reciprocals. i would make life easy and rewrite as \(5\sec^2(t)\) as a first step

OpenStudy (anonymous):

you cannot take the derivative of the denominator to solve this

OpenStudy (anonymous):

just like the derivative of \(\frac{5}{x^2}\) is not \(\frac{5}{2x}\)

OpenStudy (anonymous):

ok but all of these answers you are saying are not any of my multiple choice answers

OpenStudy (anonymous):

\[\frac{d}{dt}m(t)=\frac{d(5\sec^2 t)}{dt}=5\frac{d\sec^2 t}{d\sec t}\frac{d\sec t}{dt}=10\sec^2 t \tan x\]

OpenStudy (anonymous):

I have\[\frac{ 10\sin t }{ \cos^3 t }\]

OpenStudy (anonymous):

First of all, I have a typo lol. Secondly, \[\frac{10\sin t}{\cos^3 t}=10\frac{\sin t}{\cos t}\frac{1}{\cos^2 t}=10\tan t \sec^2 t\]

OpenStudy (anonymous):

I am getting really confused so would the answer I just gave be correct

OpenStudy (anonymous):

You can try to manipulate the answer you have.

OpenStudy (anonymous):

would \[\frac{ 10 \sin t }{ \cos t}\]

OpenStudy (anonymous):

is that a better answer

OpenStudy (anonymous):

\[\frac{10\sin t}{\cos t}\neq \frac{10 \sin t}{\cos^3 t}=10\tan t \sec^2 t\]

OpenStudy (anonymous):

um I don't have that for one of my choices

OpenStudy (anonymous):

What choices do you have?

OpenStudy (anonymous):

I have\[\frac{ -10 \sin t }{ \cos^3 t }\]

OpenStudy (anonymous):

I have\[\frac{ 10 \sin t }{ \cos^3 t }\]

OpenStudy (anonymous):

I have \[\frac{ 10 \sin t }{ \cos t }\]

OpenStudy (anonymous):

I have \[\frac{ 5 }{ 2sint \cos t }\]

OpenStudy (anonymous):

I have \[\frac{ -5 }{ 2\sin t \cos t }\]

OpenStudy (anonymous):

I have 5tant

OpenStudy (anonymous):

I have \[\frac{ -10 \sin t }{ \cos t }\]

OpenStudy (anonymous):

and I have -5tan t

OpenStudy (anonymous):

Those are all of my choices that I have

OpenStudy (anonymous):

Do you miss out a negative sign in the question?

OpenStudy (anonymous):

no negative in the question

OpenStudy (anonymous):

Can anyone help me please

OpenStudy (ash2326):

@farmergirl411 did you get the answer?

OpenStudy (anonymous):

no I am reallly stuck what I thought was the answer everyone told me it was not. I have multiple choice answers. If you scroll up a bit every post where I said I have, I have is a multiple choice answer I to choose from

OpenStudy (ash2326):

I saw your question. I'll try to help you :)

OpenStudy (anonymous):

ok thanks you so much:)

OpenStudy (ash2326):

\[m(t)=\frac{5}{\cos^2 t}\] We know that \[\frac 1 {\cos t} =\sec t\] so \[m(t)=5 \sec^2 t\] Let's take the derivative of this \[m'(t)=5\times 2 \times {\sec^{2-1} t}\times \sec t\times \tan t\] we'd get \[m'(t)=10\sec^2 t\times \tan t\] Do you get till here?

OpenStudy (anonymous):

yes

OpenStudy (ash2326):

\[m'(t)=10\times \frac 1 {\cos^2 t}\times \frac{\sin t}{\cos t}\] We get finally \[m'(t)=10\times \frac{\sin t}{\cos^3 t}\]

OpenStudy (ash2326):

I think it does match one of your option, doesn't it?

OpenStudy (anonymous):

yes it does only can I put 10 next to the sint then it will match a choice

OpenStudy (ash2326):

yes :)

OpenStudy (anonymous):

awesome thaank you sooooooooo much:)